A-Zone Media sells two models of e-readers. The budgeted price per unit for the wireless model is $193 and the budgeted price per unit for the wireless and cellular model is $418. The master budget called for sales of 10,100 wireless models and 2,550 wireless and cellular models during the current year. Actual results showed sales of 7,600 wireless models, with a price of $205 per unit, and 4,150 wireless and cellular models, with a price of $420 per unit. The standard variable cost per unit is $81 for a wireless model and $165 for a wireless and cellular model.

Required:

a. Compute the activity variance for these data. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)

b. Compute the mix and quantity variance for these data. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, or "U" for unfavorable. If there is no effect, do not select either option.)
Answer #1

Answer

a)

Activity Variance = Budegeted Sales - Actia; Sales) x Contribution Margin per unit

(10,100 - 7600 ) x (193 - 81) $ 2,80,000 Unfavorable

( 2550 - 4150) x ( 418 - 165) $ 417,450 Favorable

-----------------

697,450 Favorable

-------------------

  

b)

Mix Variance:

Flexible Budget AQ x (STD Price - STD Variable cost)   7600 x (193 - 81)   $ 851,200

4150 x ( 418-165)   $ 1,049,950

-----------------------

   $ 1,901,150

ASQ x (STD Price - STD Variable cost)

(7600 + 4150 ) x (10,100 / 12650 ) x (193-81) =    1,052,800

(7600 + 4150 ) x ( 2550 / 12650 ) x (418 -165) = 594,550 $ 1,647,350

----------------

Mix Variable    $3,548,500

------------------

Quantity Variance:

ASQ x (STD Price - STD Variable cost)

(7600 + 4150 ) x (10,100 / 12650 ) x (193-81) =    1,052,800

(7600 + 4150 ) x ( 2550 / 12650 ) x (418 -165) = 594,550

-------------------

1,647,350

Master Budget 10,100 x (193 - 81) 1,131,200

2550 x (418 165) 645,150

---------------

1,776,350

------------------

Quantity Variance 3,423,700

------------------

Cross check of Answer ( 3548500 - 3423700) = 1,24,800 Favorable

-------------------------------

If you have any query regarding the answer please as me in the comment I am here for help you. Please do not direct thumbs down just ask if you have any query. And if you like my work then please appreciates with up vote
Thank you Very much.